# BINAQUAL: A Full-Reference Objective Localization Similarity Metric for Binaural Audio

### v1

@@ 5.1 Localization Sensitivity
As shown in Fig. 2, for all elevations except ±90°, the localization similarity between test and reference signals decreases as the test azimuth moves further away from the reference. However, an increase in LS is observed between -120° to -180° and 120° to 180°, which can be attributed to front-back confusion. We also observe that the LS curve remains flat for elevations of ±90°. This can be because binaural localization primarily relies on ITD and ILD, which are minimized at extreme elevations, making azimuth localization ineffective. Instead, spectral cues dominate, but they provide poor azimuth discrimination, leading to a relatively uniform LS across azimuths.
@@
Figure 2: Localization similarity as a function of azimuth and elevation with fixed reference audio source, localized at azimuth = 0°, elevation = 0° for (a) castanets and (b) female w. reverb test signals.
Figure 3: Localization similarity scores distributed on a half sphere for various contents. The corresponding reference audio sources were localized at azimuth = 0°, elevation = 0°. A Box-Cox transform was applied to LS scores to spread out the distribution and make variations more noticeable.
@@ 5.2 Angle Interpolations
5.3 Surround Speaker Layouts
Localization similarity between samples in the surround layouts dataset–rendered for various surround speaker layouts–and reference signals (without any speaker layout) was calculated. Fig. 5 presents the aggregated mean LS scores across different speaker layouts for two scenarios: (a) source angles without elevation and (b) source angles with elevation. The 95% confidence intervals are also shown.
As shown in part (a) of Fig. 5, when the angles have no elevation, the LS scores for 5.1 and 5.1.4 are similar, as are those for 7.1 and 7.1.4. Additionally, LS scores for 7.1 and 7.1.4 are higher than those for 5.1 and 5.1.4. This suggests that increasing the number of surround channels (from 5.1 to 7.1) enhances localization similarity due to improved spatial coverage of surround speakers.
As shown in part (b) of Fig. 5, when angles have elevations, the LS scores for layouts with height channel (5.1.4, 7.1.4) are higher than those for layouts without elevation (5.1, 7.1). This confirms that elevated angles benefit more from overhead speakers, as adding height channels allows the model to capture vertical localization cues more effectively, leading to an overall increase in LS.
These results demonstrate that BINAQUAL can distinguish between simulated binaural representations of different surround speaker layouts, regardless of whether elevation is present.
Figure 5: The average localization similarity between multi-point source samples of the surround layouts dataset, rendered for different speaker layouts, and corresponding reference signals without any layouts. (a) Point source angles without elevation, (b) Point source angles with elevation.
@@ 5.4 Codec Compression
There is a strong correlation between subjective MUSHRA results and LS scores. For single-point source results, the Pearson and Spearman correlations are 0.85 and 0.84, respectively. For multiple-point source results, these correlations increase to 0.87 and 0.92, respectively, suggesting a slightly stronger monotonic relationship in the multiple-point scenario. The results indicate that the BINAQAUL model effectively distinguishes between different compression levels of binaural audio signals while remaining robust to variations in content.
@@ 5.5 General Discussion
The results presented in this study demonstrate that the BINAQUAL is effective in assessing spatial localization similarity in binaural audio across various conditions. In Sec. 5.1, we analyzed the model’s sensitivity to source localization and showed that BINAQUAL successfully detects small variations in point-source locations, as evidenced by the decrease in LS scores with increasing angular distance from the reference. However, for angles in the range of ±120° to ±180°, LS scores increased, which could be attributed to front-back confusion where binaural cues are ambiguous due to the lack of sufficient spectral differentiation. This observation aligns with psychoacoustic phenomena in binaural hearing [20]. Additionally, we found that the model is less effective in localizing pure tones due to their lack of rich interaural cues, a limitation that is consistent with human difficulties in localizing narrow-band audio sources [43]. The results also indicate that BINAQUAL maintains relatively flat LS scores at extreme elevations (±90°). This behavior is likely due to the nature of ITD and ILD cues, which are minimized at these positions. Unlike azimuthal localization, elevation perception is largely dependent on spectral cues that are harder to model robustly.
@@
Sec. 5.3 showed that BINAQUAL could differentiate effectively between 5.1-, 5.1.4-, 7.1-, and 7.1.4-rendered audio signals, showing that surround speaker layouts with more channels (and especially additional height channels) increase localization similarity. The capability to discriminate among different speaker layouts is particularly promising for applications involving complex multichannel setups in VR, AR, and cinematic audio.
@@
We addressed all five research questions outlined in Section 1, and the results demonstrated the capabilities of the BINAQUAL model. While most existing spatial localization quality metrics focus primarily on point-source localization accuracy, we took a broader approach to spatial audio evaluation. By examining the effects of angle interpolations, speaker layouts, codec compression, and content variations, our work provides deeper insights into how spatial fidelity is preserved across diverse scenarios. These aspects are crucial for advancing localization quality assessment models and ensuring objective metrics align more closely with perceptual realism. Consequently, this study not only introduces a novel localization quality assessment metric but also sets a foundation for a more holistic approach to spatial localization quality evaluation beyond traditional accuracy-based methods.
@@ 6 Conclusion
This study presented a full-reference objective localization similarity metric for binaural audio. Through a comprehensive set of validation experiments, we demonstrated the model’s capabilities in differentiating between variations in source positions, angle interpolations, surround speaker layouts, and audio degradations.
The model successfully differentiates subtle spatial variations, as evidenced by its sensitivity to small angular deviations. However, its performance is affected by well-known psychoacoustic limitations, such as front-back confusion and the reduced localization accuracy of pure tones. These findings align with established auditory localization principles, reinforcing the model’s validity. Also, BINAQUAL effectively distinguishes between different surround speaker layouts, highlighting its potential for evaluating complex binaural audio formats.
@@
Our work expands spatial audio quality evaluation beyond point-source localization accuracy by considering factors like angle interpolations, speaker layouts, compression artifacts, and content variations, offering a more comprehensive assessment of spatial fidelity in binaural audio. Future work could further refine BINAQUAL’s performance on extreme elevations and high-frequency contents to ensure even greater alignment with human spatial perception.
@@ 7 Acknowledgment
Bursty pink noise Pure tone (4 kHz) Localization based on both ITDs and ILDs
Male speech Female speech w. reverb Localizing two speech sources with different frequency ranges
Bursty white noise Castanets w. reverb Localizing an impulsive sound against a continuous bursty sound
Glockenspiel Male speech Localizing a harmonic against a vocal sound
Bursty pink noise w. reverb Female speech w. reverb Localizing two reverberant sounds
Pure tone (4 kHz) Glockenspiel Localizing two sounds with overlapping high frequencies
Bursty white noise Bursty pink noise w. reverb Localizing two sounds with broad bandwidths
Glockenspiel Castanets w. reverb Localizing two high-frequency harmonic sounds

### v2

@@ 4.4 Codec Compression Dataset
Figure 2: Localization similarity as a function of azimuth and elevation with fixed reference audio source, localized at azimuth=0°, elevation=0° for (a) Castanets and (b) Female speech w. reverb test signals.
@@ 5.1 Localization Sensitivity
As shown in Fig. 2, for all elevations except ±90°, the localization similarity between test and reference signals decreases as the test azimuth moves further away from the reference. However, an increase in LS is observed between -120° to -180° and 120° to 180°, which can be attributed to front-back confusion. For elevations of ±90°, the LS curve remains relatively flat across azimuths. This is expected, as at ±90° elevation, corresponding to directly above or below the listener, azimuth resolution effectively collapses. In other words, in Cartesian coordinates, all azimuths converge to a single point, making azimuth distinctions meaningless. Additionally, binaural localization cues such as ITD and ILD become minimal at these elevations, and although spectral cues are still present, they offer limited azimuthal discrimination, further contributing to a uniform LS profile.
@@
Figure 3: Localization similarity scores distributed on a half sphere for various contents. The corresponding reference audio sources were localized at azimuth = 0°, elevation = 0°. Box-Cox transform was applied to LS scores to spread out the distribution and make variations more noticeable.
@@ 5.2 Angle Interpolations
5.3 Surround Loudspeaker Layouts
Localization similarity between samples in the surround layouts dataset, rendered for various surround loudspeaker layouts, and reference signals (without any loudspeaker layout) was calculated. Fig. 5 presents the aggregated mean LS scores across different loudspeaker layouts for two scenarios: (a) source angles without elevation and (b) source angles with elevation. The mean LS scores were computed across 8 different audio samples per layout. For each configuration, the 95% confidence interval was also calculated using the standard error of the mean across the 8 samples.
As shown in part (a) of Fig. 5, when the angles have no elevation, the LS scores for 5.1 and 5.1.4 are similar, as are those for 7.1 and 7.1.4. Additionally, LS scores for 7.1 and 7.1.4 are higher than those for 5.1 and 5.1.4. This suggests that increasing the number of surround channels enhances localization similarity due to improved spatial coverage of surround loudspeakers.
As shown in part (b) of Fig. 5, when angles have elevations, the LS scores for layouts with height channel (5.1.4, 7.1.4) are higher than those for layouts without elevation (5.1, 7.1). This confirms that elevated angles benefit more from overhead loudspeakers, as adding height channels allows the model to capture vertical localization cues more effectively, leading to an overall increase in LS.
Statistical analysis using non-parametric tests further supports these findings. As shown in Table A6, Mann–Whitney U tests reveal that increasing the number of surround channels leads to significantly higher LS scores, as evidenced by comparisons between 5.1 vs. 7.1 and 5.1.4 vs. 7.1.4 (both with $p=0.0002$). Similarly, adding height channels significantly improves localization similarity for both 5.1 vs. 5.1.4 ($p=0.0499$) and 7.1 vs. 7.1.4 ($p=0.0003$). Kruskal–Wallis tests across all layouts further confirm overall significant differences due to layout configuration. These results demonstrate that BINAQUAL can distinguish between simulated binaural representations of different surround loudspeaker layouts, regardless of whether elevation is present.
Figure 5: The average localization similarity between multi-point source samples of the surround layouts dataset, rendered for different loudspeaker layouts, and corresponding reference signals without any layouts. (a) Point source angles without elevation, (b) Point source angles with elevation.
@@ 5.4 Codec Compression
There is a strong correlation between subjective MUSHRA results and LS scores. For single-point source results, the Pearson and Spearman correlations are 0.85 and 0.84, respectively. For multiple-point source results, these correlations increase to 0.87 and 0.92, respectively, suggesting a slightly stronger monotonic relationship in the multiple-point scenario.
Additionally, we computed the correlation between AMBIQUAL localization accuracy scores and BINAQUAL LS scores. For single-point sources, the Pearson and Spearman correlation coefficients are 0.84 and 0.80, respectively. For multiple-point sources, the correlations increase to 0.93 (Pearson) and 0.96 (Spearman).
The results indicate that the BINAQAUL model effectively distinguishes between different compression levels of binaural audio signals while remaining robust to variations in content.
@@ 5.5 General Discussion
The results presented in this study demonstrate that the BINAQUAL is effective in assessing spatial localization similarity in binaural audio across various conditions. In Sec. 5.1, we analyzed the model’s sensitivity to source localization and showed that BINAQUAL successfully detects small variations in point-source locations, as evidenced by the decrease in LS scores with increasing angular distance from the reference. However, for angles in the range of ±120° to ±180°, LS scores increased, which could be attributed to front-back confusion where binaural cues are ambiguous due to the lack of sufficient spectral differentiation. This observation aligns with psychoacoustic phenomena in binaural hearing [21]. Additionally, we found that the model is less effective in localizing pure tones due to their lack of rich interaural cues, a limitation that is consistent with human difficulties in localizing narrow-band audio sources [51]. The results also indicate that BINAQUAL maintains relatively flat LS scores at extreme elevations (±90°). This behavior is expected, as azimuth resolution effectively collapses at ±90° elevation. Also, elevation perception relies more heavily on spectral cues, which are typically more variable and harder to model robustly, further contributing to the observed uniformity in LS scores.
@@
Sec. 5.3 showed that BINAQUAL could differentiate effectively between 5.1-, 5.1.4-, 7.1-, and 7.1.4-rendered audio signals, showing that surround loudspeaker layouts with more channels (and especially additional height channels) increase localization similarity. The capability to discriminate among different loudspeaker layouts is particularly promising for applications involving complex multichannel setups in VR, AR, and cinematic audio.
@@
While BINAQUAL effectively generalizes across content, source configurations, and degradations, some limitations remain. Its performance declines in conditions where binaural cues are inherently ambiguous or diminished, such as at extreme elevations, with pure tone stimuli, or in cases of front-back confusion. These are also well-known challenges for human listeners, which reinforces the perceptual validity of the metric. We should note that our goal is not to outperform human hearing, but rather to develop a metric that closely mimics how humans perceive spatial audio quality. Future work could enhance BINAQUAL by integrating additional spectral cues or simulating listener head movements, strategies that humans naturally use to resolve spatial ambiguities. Such enhancements aim to further align the metric with perceptual realism and improve its applicability to dynamic or immersive spatial audio scenarios.
We addressed all five research questions outlined in Section 1, and the results demonstrated the capabilities of the BINAQUAL model. While most existing spatial localization quality metrics focus primarily on point-source localization accuracy, we took a broader approach to spatial audio evaluation. By examining the effects of angle interpolations, loudspeaker layouts, codec compression, and content variations, our work provides deeper insights into how spatial fidelity is preserved across diverse scenarios. These aspects are crucial for advancing localization quality assessment models and ensuring objective metrics align more closely with perceptual realism. Consequently, this study not only introduces a novel localization quality assessment metric but also sets a foundation for a more holistic approach to spatial localization quality evaluation beyond traditional accuracy-based methods.
@@ 6 CONCLUSION
This study presented a full-reference objective localization similarity metric for binaural audio. Through a comprehensive set of validation experiments, we demonstrated the model’s capabilities in differentiating between variations in source positions, angle interpolations, surround loudspeaker layouts, and audio degradations.
The model successfully differentiates subtle spatial variations, as evidenced by its sensitivity to small angular deviations. However, its performance is affected by well-known psychoacoustic limitations, such as front-back confusion and the reduced localization accuracy of pure tones. These findings align with established auditory localization principles, reinforcing the model’s validity. Also, BINAQUAL effectively distinguishes between different surround loudspeaker layouts, highlighting its potential for evaluating complex binaural audio formats.
@@
Our work expands spatial audio quality evaluation beyond point-source localization accuracy by considering factors like angle interpolations, loudspeaker layouts, compression artifacts, and content variations, offering a more comprehensive assessment of spatial fidelity in binaural audio. Future work could further refine BINAQUAL’s performance to ensure even greater alignment with human spatial perception.
@@ 7 Acknowledgment
Male speech Female speech w. reverb Localizing two vocal sources with different frequency ranges
Castanets Male speech Localizing a percussive natural sound against a vocal source
Bursty pink noise Female speech Localizing a synthetic broadband sound against a vocal source
Bursty white noise Male speech Localizing a synthetic broadband sound against a vocal source
Bursty pink noise w. reverb Female speech w. reverb Localizing two reverberant sounds with different characteristics
Bursty white noise Bursty pink noise w. reverb Localizing two synthetic broadband sounds with distinct temporal properties
Bursty pink noise Castanets Localizing a synthetic sound against a natural percussive sound
Bursty white noise Castanets w. reverb Localizing a synthetic sound against a reverberant percussive natural sound
@@
Table A6: Statistical comparison between different surround loudspeaker layouts using Mann–Whitney and Kruskal–Wallis tests.
Effect of increasing the layout size
Comparison Test p-value
5.1 vs 7.1 Mann–Whitney U 0.0002
5.1.4 vs 7.1.4 Mann–Whitney U 0.0002
All layouts Kruskal–Wallis 0.0000
Effect of adding more height channels
Comparison Test p-value
5.1 vs 5.1.4 Mann–Whitney U 0.0499
7.1 vs 7.1.4 Mann–Whitney U 0.0003
All layouts Kruskal–Wallis 0.0008